OpenStudy (anonymous):

medal for the correct answer! Let f(x) = 9x - 2 and g(x) = -x + 3. Find f(g(x)). -9x - 2 -9x + 5 -9x + 25 -9x + 27

OpenStudy (anonymous):

\[f \left( g \left( x \right) \right)=f \left( -x+3 \right)=9\left( -x+3 \right)-2\] solve and get the answer.